# Classical simulation of non-Gaussian bosonic circuits

### v1

@@ 5 Superpositions of Gaussian states and measurements
$\left\{c_{j},\Delta(\psi_{j})\right\}_{j\in[\chi]}$ (144)
of coefficients and descriptions of the individual terms in the superposition. We describe how to compute or estimate the probability density functions when applying a homodyne measurement to a superposition of Gaussian states specified by this data. We also explain how to update descriptions of such superpositions after a Gaussian unitary or measurement is applied. This relies on the routines outlined in Section 4 for tracking relative phases in linear optics, as well as (in the case of measurements) on the norm estimation algorithms laid out in Sections 5.1 and 5.2. For a more detailed explanation of this procedure, we refer readers to [21, Sections 4.1 and 4.3] where analogous subroutines are discussed in the context of fermionic linear optics.
@@ 5.2 A fast norm estimation algorithm
In pioneering work [20], Bravyi and Gosset proposed an algorithm for estimating the norm of a superposition of stabilizer states that takes time linear in the number of terms $\chi$ in the superposition. In Ref. [33] the same authors proposed a similar algorithm to estimate the norm of superpositions of fermionic Gaussian states that also takes time linear in the number of terms $\chi$ in the superposition. These algorithms correspond to an improvement over the naive norm computation which takes time $O(\chi^{2})$.
@@
$\langle\Psi^{\prime},H\Psi^{\prime}\rangle\leq E\qquad\textrm{ where }\qquad\Psi^{\prime}=\Psi/\|\Psi\|\ .$ (149)
@@
$R:=\sqrt{\frac{E}{\varepsilon}}\qquad\textrm{ and }\qquad L:=\left\lceil\frac{E}{4\pi p_{f}\varepsilon^{3}}\right\rceil\ .$ (150)
@@
[…]
@@
$p(\alpha)$ $=\frac{1}{\pi}\frac{|\langle\alpha,\Psi\rangle|^{2}}{\|\Psi\|^{2}}\qquad\textrm{ for }\qquad\alpha\in\mathbb{C}\ .$ (154)
Then $p$ is a probability density function (the one associated with the outcome distribution when applying the heterodyne measurement to the normalized state Ψ/$\Psi/\|\Psi\|$), and we have
@@
[…]
@@
[…]
@@
[…]
@@
$R$ $\geq\sqrt{\frac{\langle\Psi^{\prime},H\Psi^{\prime}\rangle}{\varepsilon}}\qquad\textrm{ and }\qquad L\geq\frac{R^{2}}{4\pi p_{f}\varepsilon^{2}}\ .$ (177)
@@
$R:=\sqrt{\frac{E}{\varepsilon}}\qquad\textrm{ and }\qquad L:=\left\lceil\frac{E}{4\pi p_{f}\varepsilon^{3}}\right\rceil$ (178)
@@
1:$R\leftarrow\sqrt{\frac{E}{\varepsilon}}$
2:$L\leftarrow\big{\lceil}\frac{E}{4\pi p_{f}\varepsilon^{3}}\big{\rceil}$
@@
Figure 11: The algorithm $\mathsf{fastnorm}_{E}$ takes as input a description of a $n$-mode superposition $\Psi$ of $\chi$ Gaussian states and parameters $\varepsilon>0,p_{f}\in(0,1)$, and it returns an estimate of the norm squared $\|\Psi\|^{2}$ up to multiplicative error $\varepsilon$ with success probability $1-p_{f}$. It requires an upper bound $E$ on the energy of the normalized state Ψ/$\Psi/\|\Psi\|$.
It is clear that this can be turned into a probabilistic algorithm for estimating the quantity $\|\Psi\|^{2}$ up to a multiplicative error. We give such an algorithm in Fig. 11 and we establish the associated claim in Theorem 5.3. Theorem 5.3 is an immediate consequence of Lemma 5.2.
@@
The algorithm $\mathsf{fastnorm}_{E}$ given in Fig. 11 achieves the following, given as input a tuple $\left\{c_{j},\Delta(\psi_{j})\right\}_{j\in[\chi]}$ specifying a superposition $|\Psi\rangle=\sum_{j=1}^{\chi}c_{j}|\psi_{j}\rangle\in\mathcal{H}_{n}$ of Gaussian states $\psi_{j},j\in[\chi]$ where
$\langle\Psi^{\prime},H\Psi^{\prime}\rangle\leq E\qquad\textrm{ with }\qquad\Psi^{\prime}=\Psi/\|\Psi\|\ ,$ (179)
@@
$\langle\Psi^{\prime},H\Psi^{\prime}\rangle\leq E\qquad\textrm{ where }\qquad\Psi^{\prime}=\Psi/\|\Psi\|\ .$ (182)
@@
in Line 7. Then, in Line 147 we have
@@
Because of its central relevance for the complexity of our algorithm, let us briefly comment on how the energy (i.e., mean photon number) of the states occuring during the computation can be bounded. The corresponding bound depends on the mean photon number $N$ of the initial superposition, as well as the unitaries applied. Indeed, one can use that these unitaries correspond to energy-limited channels.
Recall we use the Hamiltonian given in Eq. 29. A channel $\mathcal{N}$ is said to be energy-limited if it maps energy bounded states to energy bounded states [34, 35]. Gaussian channels including finite displacements, phase shifters, beamsplitters and finite squeezing are energy-limited. Concretely, we have
$\operatorname{tr}(HD(\alpha)^{\dagger}\rho D(\alpha))$ $=\operatorname{tr}(H\rho)+d(\alpha)^{T}d(\alpha)+2s^{T}d(\alpha)$ for $\qquad\text{for all}\qquad\alpha\in\mathbb{C}^{n}\ ,$ (187)
where $s\in\mathbb{R}^{2n}$ is the displacement of the state $\rho$, and
$\operatorname{tr}(HP_{j}(\phi)^{\dagger}\rho P_{j}(\phi))$ $=\operatorname{tr}(H\rho)$ for $\qquad\text{for all}\qquad\phi\in\mathbb{R}\ ,$ (188)
$\operatorname{tr}(HB_{j,k}(\phi)^{\dagger}\rho B_{j,k}(\phi))$ $=\operatorname{tr}(H\rho)$ for $\qquad\text{for all}\qquad\omega\in\mathbb{R}\ ,$ (189)
$\operatorname{tr}(HS_{j}(z)^{\dagger}\rho S_{j}(z))$ $\leq e^{2z}\operatorname{tr}(H\rho)$ for $\qquad\text{for all}\qquad z>1\ ,$ (190)
@@
[…]
@@ 5.3 Measurement of a superposition
$p_{\Psi}(\alpha)$ $=\frac{1}{\pi^{k}}\|\Pi_{\alpha}\Psi\|^{2}\qquad\textrm{ where }\qquad\Pi_{\alpha}=|\alpha\rangle\langle\alpha|\otimes I^{\otimes n-k}$ (196)
@@
$\Pi_{\alpha}\Psi$ $=\sum_{j=1}^{\chi}c_{j}\Pi_{\alpha}\psi_{j}$ (197)
$=\sum_{j=1}^{\chi}c^{\prime}_{j}(\alpha)\psi^{\prime}_{j}(\alpha)\qquad\textrm{ where }\qquad\begin{matrix}c^{\prime}_{j}(\alpha)&=&\kern-20.00003ptc_{j}\|\Pi_{\alpha}\psi_{j}\|\\ \psi^{\prime}_{j}(\alpha)&=&\Pi_{\alpha}\psi_{j}/\|\Pi_{\alpha}\psi_{j}\|\end{matrix}$ (198)
@@
The algorithm $\mathsf{measureprobapproximate}_{E}$ corresponds to running the algorithm $\mathsf{fastnorm}_{E}$ with input $\{c_{j}^{\prime}(\alpha),\Delta(\psi_{j}^{\prime}(\alpha))\}_{j\in[\chi]}$, $\varepsilon$ and $p_{f}$. By Theorem 5.3, dividing by $\pi^{k}$ this gives a value $p$ satisfying the claim. As for the exact algorithm, this requires $\chi$ uses of the algorithms $\mathsf{prob}$ and $\mathsf{postmeasure}$. Then, the runtime of $\mathsf{measureprobexact}$ is
@@
for any $R>0$ by the same reasoning as in the proof of Eq. (156). That is, most of the probability mass of the distribution $p_{\Psi}$ is contained within a ball of radius $R$ dictated by the energy of $\Psi$. This means that the computational problem of determining $p_{\Psi}(\alpha)$ will typically only be of interest for $\alpha\in\mathbb{C}^{k}$ with small norm. Furthermore, for such “typical” measurement outcomes $\alpha\in\mathbb{C}^{k}$, the post-measurement state $\Psi^{\prime}_{\alpha}$ has small norm. Indeed, since the heterodyne measurement channel $\mathcal{E}$ is energy-limited (see Eqs. 191 and 192) we have
[…]
By Markov’s inequality, this implies that the distribution $p_{\Psi}$ has support concentrated on $\alpha\in B_{R}(0)\subset\mathbb{C}^{k}$ with the property that the associated post-measuremen state $\langle\Psi^{\prime}_{\alpha},H\Psi^{\prime}_{\alpha}\rangle$ has bounded energy (compared to that of $\Psi$).
We can summarize these observation by the following lemma. It shows that the estimation of $p_{\Psi}(\alpha)$ works efficiently whenever $\alpha$ belongs to a “typical” set $\mathcal{T}_{\delta}$. That is, we have:
@@
$\Pr\left[|\alpha\|^{2}\geq R^{2}\right]$ $\leq\delta/2$ (211)
@@
$\mathcal{T}_{\delta}$ :=$:=\left\{\alpha\in B_{R}(0)\ |\ \langle\Psi_{\alpha}^{\prime},H\Psi_{\alpha}^{\prime}\rangle\leq\tilde{E}\right\}\subset\mathbb{C}^{k}\ ,$ (212)
@@
$\Pr\left[\alpha\in\mathcal{T}_{\delta}\right]\geq 1-\delta$ (213)
@@ Acknowledgements
$\hat{d}_{2j-1}(\alpha)&=\sqrt{2}\operatorname{Re}(\alpha_{j})\\ \hat{d}_{2j}(\alpha)&=\sqrt{2}\operatorname{Im}(\alpha_{j})\end{aligned}\end{gathered}\qquad\text{for}\qquad j\in[n]\ .$ (216)
@@
$\alpha_{j}$ $=\frac{1}{\sqrt{2}}(d_{2j-1}+id_{2j})\qquad\textrm{ for }\qquad j\in[n]\ ,$ (217)
@@
$d_{j}$ $=\langle\psi,R_{j}\psi\rangle\qquad\text{for}\qquad j\in[2n]$ (222)
@@
$\lambda_{j}$ $\geq 4\qquad\textrm{ for all }j\in[n]\ ,$ (227)
@@
$=h(\lambda_{j},\lambda_{k})\prod_{\ell\in\{1,\ldots,n\}\backslash\{j,k\}}\lambda_{\ell}\qquad\textrm{ where }\qquad h(x,y):=xy\ .$ (234)
It is straightforward to check that for any $\varepsilon\geq 8$
@@
$=(\Lambda-4(n-1))4^{n-1}$ (237)
$=\left(2\sum_{j=1}^{2n}\operatorname{Var}_{\psi}(R_{j})-2n+4\right)4^{n-1}$ (238)
@@
$\operatorname{tr}\left(D(\alpha)\rho(\operatorname{\Gamma}_{1},d_{1})\rho(\operatorname{\Gamma}_{2},d_{2})\rho(\operatorname{\Gamma}_{3},0)\right)\qquad\text{for}\qquad\alpha\in\mathbb{C}^{n}\ .$ (242)
@@
$\operatorname{tr}\left(D(\alpha)\rho(\operatorname{\Gamma}_{1},d_{1})\rho(\operatorname{\Gamma}_{2},d_{2})\rho(\operatorname{\Gamma}_{3},0)\right)$ $=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{2n}}d^{2n}\xi\,\overline{\chi_{\rho(\operatorname{\Gamma}_{1},d_{1})D(-\alpha)}(\xi)}\chi_{\rho(\operatorname{\Gamma}_{2},d_{2})\rho(\operatorname{\Gamma}_{3},0)}(\xi)\ .$ (243)
@@
$\operatorname{tr}(D(\xi))$ $=\delta(\xi)\qquad\text{ for any }\qquad\xi\in\mathbb{R}^{2n}\ ,$ (253)
$\frac{1}{2\pi}\int_{\mathbb{R}^{2n}}d^{2n}x\,f(x)\delta(x-a)$ $=f(a)\qquad\text{ for any }\qquad a\in\mathbb{R}^{2n}\ ,$ (254)
@@
Replacing $M$ and $y$ respectively by Eqs. 265 and 266 and additional algebraic manipulation gives the claim. ∎
Inserting Eqs. 249 and 257 into Eq. 243 gives
@@
Replacing $M$ and $y$ respectively by Eqs. 273 and 274 and additional algebraic manipulation gives
@@
with $\Gamma_{5},W_{1},W_{2},W_{3},W_{4},W_{5}$ defined in Eqs. 80, 81, 82, 83, 84 and 85.

### v2

@@ 4 Superpositions of Gaussian states and measurements
$\left\{c_{j},\Delta(\psi_{j})\right\}_{j\in[\chi]}\$(149)
of coefficients and descriptions of the individual terms in the superposition. We describe how to compute or estimate the probability density functions when applying a homodyne measurement to a superposition of Gaussian states specified by this data. We also explain how to update descriptions of such superpositions after a Gaussian unitary or measurement is applied. This relies on the routines outlined in Section 3 for tracking relative phases in linear optics, as well as (in the case of measurements) on the norm estimation algorithms laid out in Sections 4.1 and 4.2. For a more detailed explanation of this procedure, we refer readers to [dias2023classical, Sections 4.1 and 4.3] where analogous subroutines are discussed in the context of fermionic linear optics.
@@ 4.2 A fast norm estimation algorithm
In pioneering work [bravyiImprovedClassicalSimulation2016], Bravyi and Gosset proposed an algorithm for estimating the norm of a superposition of stabilizer states that takes time linear in the number of terms $\chi$ in the superposition. In Ref. [bravyiComplexityQuantumImpurity2017a] the same authors proposed a similar algorithm to estimate the norm of superpositions of fermionic Gaussian states that also takes time linear in the number of terms $\chi$ in the superposition. These algorithms correspond to an improvement over the naive norm computation which takes time $O(\chi^{2})$.
@@
$\langle\Psi^{\prime},H\Psi^{\prime}\rangle\leq E\quad\textrm{where}\quad\Psi^{\prime}=\Psi/\|\Psi\|\ .$ (154)
@@
$R:=\sqrt{\frac{E}{\varepsilon}}\quad\textrm{ and }\quad L:=\left\lceil\frac{E}{4\pi p_{f}\varepsilon^{3}}\right\rceil\ .$ (155)
@@
$\langle\Psi,H\Psi\rangle$ $=\frac{1}{\pi}\int_{\mathbb{C}}2|\alpha|^{2}|\langle\alpha,\Psi\rangle|^{2}d\alpha\ ,$ (158)
@@
$p(\alpha)$ $=\frac{1}{\pi}\frac{|\langle\alpha,\Psi\rangle|^{2}}{\|\Psi\|^{2}}\quad\textrm{for}\quad\alpha\in\mathbb{C}\ .$ (159)
Then $p$ is a probability density function (the one associated with the outcome distribution when applying the heterodyne measurement to the normalized state $\Psi/\|\Psi\|$), and we have
@@
$\mathbb{E}\left[X\right]$ $=\frac{1}{\pi R^{2}}\int_{B_{R}(0)}R^{2}|\langle\alpha,\Psi\rangle|^{2}d\alpha$ (163)
$=\frac{1}{\pi}\int_{B_{R}(0)}|\langle\alpha,\Psi\rangle|^{2}d\alpha$ (164)
$=\|\Psi\|^{2}\int_{B_{R}(0)}p(\alpha)d\alpha\ ,$ (165)
@@
$\mathbb{E}[X^{2}]$ $=\frac{1}{\pi R^{2}}\int_{B_{R}(0)}R^{4}|\langle\alpha,\Psi\rangle|^{4}d\alpha$ (167)
$=\pi R^{2}\|\Psi\|^{4}\int_{B_{R}(0)}p(\alpha)^{2}d\alpha$ (168)
@@
where we used that $\int_{B_{R}(0)}p(\alpha)^{2}d\alpha\leq\int_{B_{R}(0)}p(\alpha)d\alpha\leq 1$. Thus
@@
$R$ $\geq\sqrt{\frac{\langle\Psi^{\prime},H\Psi^{\prime}\rangle}{\varepsilon}}\quad\textrm{and}\quad L\geq\frac{R^{2}}{4\pi p_{f}\varepsilon^{2}}\ .$ (182)
@@
$R:=\sqrt{\frac{E}{\varepsilon}}\quad\textrm{and}\quad L:=\left\lceil\frac{E}{4\pi p_{f}\varepsilon^{3}}\right\rceil$ (183)
@@
It is clear that this can be turned into a probabilistic algorithm $\mathsf{fastnorm}_{E}$ for estimating the quantity $\|\Psi\|^{2}$ up to a multiplicative error. The algorithm $\mathsf{fastnorm}_{E}$ takes as input a description of a $n$-mode superposition $\Psi$ of $\chi$ Gaussian states and parameters $\varepsilon>0,p_{f}\in(0,1)$, and it returns an estimate of the norm squared $\|\Psi\|^{2}$ up to multiplicative error $\varepsilon$ with success probability $1-p_{f}$. It requires an upper bound $E$ on the energy of the normalized state $\Psi/\|\Psi\|$. We establish the associated claim in Theorem 4.3. Theorem 4.3 is an immediate consequence of Lemma 4.2.
@@
[…]
@@
The algorithm $\mathsf{fastnorm}_{E}$ given in Section 4.2 achieves the following, given as input a tuple $\left\{c_{j},\Delta(\psi_{j})\right\}_{j\in[\chi]}$ specifying a linear combination $|\Psi\rangle=\sum_{j=1}^{\chi}c_{j}|\psi_{j}\rangle\in\mathcal{H}_{n}$ of Gaussian states $\psi_{j},j\in[\chi]$ where
$\langle\Psi^{\prime},H\Psi^{\prime}\rangle\leq E\quad\textrm{with}\quad\Psi^{\prime}=\Psi/\|\Psi\|\ ,$ (184)
@@
Note added: The procedure in Threorem 4.3 can be modified by using a median-of-means argument as done in [bravyiSimulationQuantumCircuits2019a, Section 4.1] (see also [reardon-smithImprovedSimulationQuantum2023, Lemma 10]). This gives an improved runtime of $O\left(\chi n^{3}E\varepsilon^{-3}\log(1/p_{f})\right)$.
@@
$\langle\Psi^{\prime},H\Psi^{\prime}\rangle\leq E\quad\textrm{where}\quad\Psi^{\prime}=\Psi/\|\Psi\|\ .$ (187)
@@
in Line 7. Then, in Line 152 we have
@@
Because of its central relevance to the complexity of our algorithm, let us briefly comment on how the energy (i.e., mean photon number) of the states occurring during the computation can be bounded. The corresponding bound depends on the mean photon number $N$ of the initial superposition, as well as the unitaries applied. Indeed, one can use that these unitaries correspond to energy-limited channels.
Recall we use the Hamiltonian given in Eq. 32. A channel $\mathcal{N}$ is said to be energy-limited if it maps energy bounded states to energy bounded states [winter2017energyconstrained, shirokovExtensionQuantumChannels2020]. Gaussian channels including finite displacements, phase shifters, beamsplitters and finite squeezing are energy-limited. Concretely, we have
$\operatorname{tr}(HD(\alpha)^{\dagger}\rho D(\alpha))$ $=\operatorname{tr}(H\rho)+d(\alpha)^{T}d(\alpha)+2s^{T}d(\alpha)$
for all $\alpha\in\mathbb{C}^{n}$, where $s\in\mathbb{R}^{2n}$ is the displacement of the state $\rho$, and
$\operatorname{tr}(HP_{j}(\phi)^{\dagger}\rho P_{j}(\phi))$ $=\operatorname{tr}(H\rho)$ $\quad\text{for all}\quad\phi\in\mathbb{R}\ ,$
$\operatorname{tr}(HB_{j,k}(\phi)^{\dagger}\rho B_{j,k}(\phi))$ $=\operatorname{tr}(H\rho)$ $\quad\text{for all}\quad\omega\in\mathbb{R}\ ,$
$\operatorname{tr}(HS_{j}(z)^{\dagger}\rho S_{j}(z))$ $\leq e^{2z}\operatorname{tr}(H\rho)$ $\quad\text{for all}\quad z>1\ ,$
@@
$\mathcal{E}(\rho)$ $=\frac{1}{\pi}\int_{\mathbb{C}}|\alpha\rangle\langle\alpha|\rho|\alpha\rangle\langle\alpha|d\alpha\ .$ (192)
@@ 4.3 Measurement of a superposition
$p_{\Psi}(\alpha)$ $=\frac{1}{\pi^{k}}\|\Pi_{\alpha}\Psi\|^{2}\quad\textrm{where}\quad\Pi_{\alpha}=|\alpha\rangle\langle\alpha|\otimes I^{\otimes n-k}\$(197)
@@
$\Pi_{\alpha}\Psi$ $=\sum_{j=1}^{\chi}c_{j}\Pi_{\alpha}\psi_{j}=\sum_{j=1}^{\chi}c^{\prime}_{j}(\alpha)\psi^{\prime}_{j}(\alpha)$ (198)
where
$c^{\prime}_{j}(\alpha)$ $=c_{j}\|\Pi_{\alpha}\psi_{j}\|$ (199)
$\psi^{\prime}_{j}(\alpha)$ $=\Pi_{\alpha}\psi_{j}/\|\Pi_{\alpha}\psi_{j}\|$ (200)
@@
The algorithm $\mathsf{measureprobapproximate}_{E}$ corresponds to running the algorithm $\mathsf{fastnorm}_{E}$ with input $\{c_{j}^{\prime}(\alpha),\Delta(\psi_{j}^{\prime}(\alpha))\}_{j\in[\chi]}$, $\varepsilon$ and $p_{f}$. By Theorem 4.3, dividing by $\pi^{k}$ gives a value $p$ satisfying the claim. As for the exact algorithm, this requires $\chi$ uses of the algorithms $\mathsf{prob}$ and $\mathsf{postmeasure}$. Then, the runtime of $\mathsf{measureprobexact}$ is
@@
for any $R>0$ by the same reasoning as in the proof of Eq. (161). That is, most of the probability mass of the distribution $p_{\Psi}$ is contained within a ball of radius $R$ dictated by the energy of $\Psi$. This means that the computational problem of determining $p_{\Psi}(\alpha)$ will typically only be of interest for $\alpha\in\mathbb{C}^{k}$ with small norm. Furthermore, for such “typical” measurement outcomes $\alpha\in\mathbb{C}^{k}$, the post-measurement state $\Psi^{\prime}_{\alpha}$ has small norm. Indeed, since the heterodyne measurement channel $\mathcal{E}$ is energy-limited (see Eqs. 192 and 193) we have
$\int_{\mathbb{C}^{k}}p_{\Psi}(\alpha)\langle\Psi^{\prime}_{\alpha},H\Psi^{\prime}_{\alpha}\rangle d\alpha$ $=\langle\Psi,H\Psi\rangle+1\ .$ (208)
By Markov’s inequality, this implies that the distribution $p_{\Psi}$ has support concentrated on $\alpha\in B_{R}(0)\subset\mathbb{C}^{k}$ with the property that the associated post-measurement state $\langle\Psi^{\prime}_{\alpha},H\Psi^{\prime}_{\alpha}\rangle$ has bounded energy (compared to that of $\Psi$).
We can summarize these observations by the following lemma. It shows that the estimation of $p_{\Psi}(\alpha)$ works efficiently whenever $\alpha$ belongs to a “typical” set $\mathcal{T}_{\delta}$. That is, we have:
@@
$\Pr\left[|\alpha\|^{2}\geq R^{2}\right]$ $\leq\delta/2\$(213)
@@
$\mathcal{T}_{\delta}$ $:=\left\{\alpha\in B_{R}(0)\ |\ \langle\Psi_{\alpha}^{\prime},H\Psi_{\alpha}^{\prime}\rangle\leq\tilde{E}\right\}\subset\mathbb{C}^{k}\ ,$ (214)
@@
$\Pr\left[\alpha\in\mathcal{T}_{\delta}\right]\geq 1-\delta\$(215)
@@
Note added: After posting our work to the arXiv, we were made aware of concurrent independent work by Hahn et al. [hahn2024classical], which also establishes strong simulation algorithms for non-Gaussian operations along similar lines.
@@ Acknowledgements
$\begin{gathered}\begin{aligned} \hat{d}_{2j-1}(\alpha)&=\sqrt{2}\operatorname{Re}(\alpha_{j})\\ \hat{d}_{2j}(\alpha)&=\sqrt{2}\operatorname{Im}(\alpha_{j})\end{aligned}\end{gathered}\quad\text{for}\quad j\in[n]\ .$ (218)
@@
$\alpha_{j}$ $=\frac{1}{\sqrt{2}}(d_{2j-1}+id_{2j})\quad\textrm{for}\quad j\in[n]\ ,$ (219)
@@
$d_{j}$ $=\langle\psi,R_{j}\psi\rangle\quad\text{for}\quad j\in[2n]$ (224)
@@
$\lambda_{j}$ $\geq 4\quad\textrm{ for all }j\in[n]\ ,$ (229)
@@
[…]
where $h(x,y):=xy$. It is straightforward to check that for any $\varepsilon\geq 8$
@@
$\qquad\qquad=(\Lambda-4(n-1))4^{n-1}$ (239)
$\qquad\qquad=\left(2\sum_{j=1}^{2n}\operatorname{Var}_{\psi}(R_{j})-2n+4\right)4^{n-1}$ (240)
@@
$\operatorname{tr}\left(D(\alpha)\rho(\operatorname{\Gamma}_{1},d_{1})\rho(\operatorname{\Gamma}_{2},d_{2})\rho(\operatorname{\Gamma}_{3},0)\right)\quad\text{for}\quad\alpha\in\mathbb{C}^{n}\ .$ (244)
@@
$=\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{2n}}d^{2n}\xi\,\overline{\chi_{\rho(\operatorname{\Gamma}_{1},d_{1})D(-\alpha)}(\xi)}\chi_{\rho(\operatorname{\Gamma}_{2},d_{2})\rho(\operatorname{\Gamma}_{3},0)}(\xi)\ .$ (246)
@@
$\operatorname{tr}(D(\xi))$ $=\delta(\xi)\quad\text{for any}\quad\xi\in\mathbb{R}^{2n}\ ,$ (256)
$\frac{1}{2\pi}\int_{\mathbb{R}^{2n}}d^{2n}x\,f(x)\delta(x-a)$ $=f(a)\quad\text{for any}\quad a\in\mathbb{R}^{2n}\ ,$ (257)
@@
Replacing $M$ and $y$ respectively by Eqs. 268 and 269 and additional algebraic manipulation gives the claim. ∎
Inserting Eqs. 252 and 260 into Eq. 245 gives
@@
Replacing $M$ and $y$ respectively by Eqs. 276 and 277 and additional algebraic manipulation gives
@@
with $\Gamma_{5},W_{1},W_{2},W_{3},W_{4},W_{5}$ defined in Eqs. 85, 86, 87, 88, 89 and 90.